Determine if the ordered pair a solution of the inequality ?
(3,12); 5y-6x>3
Help me with this please.

Answer by nabla(475)

The ordered pair means x=3 and y=12. Put those two values into the inequality and see if it is true.
5(12)-6(3)=60-18=42. Is 42>3? Certainly.